OpenStudy (anonymous):

The minimum and maximum temperature for a day in Cupcake Town can be modeled by the equation below: 3|x − 8| + 15 = 18 What are the minimum and maximum temperatures for this day? x = 1, x = 17 x = −7, x = 9 x = 7 x = 9 No solution

OpenStudy (nurali):

3|x − 8| + 15 = 18 |x − 8|=1 x-8=1 or x-8=-1 x=9 or x=-7